# HANDI: Hand-Centric Text-and-Image Conditioned Video Generation

### v4

@@ Abstract
Despite the recent strides in video generation, state-of-the-art methods still struggle with elements of visual detail. One particularly challenging case is the class of videos in which the intricate motion of the hand coupled with a mostly stable and otherwise distracting environment is necessary to convey the execution of some complex action and its effects. To address these challenges, we introduce a new method for video generation that focuses on hand-centric actions. Our diffusion-based method incorporates two distinct innovations. First, we propose an automatic method to generate the motion area—the region in the video in which the detailed activities occur—guided by both the visual context and the action text prompt, rather than assuming this region can be provided manually as is now commonplace. Second, we introduce a critical Hand Refinement Loss to guide the diffusion model to focus on smooth and consistent hand poses. We evaluate our method on challenging augmented datasets based on EpicKitchens and Ego4D, demonstrating significant improvements over state-of-the-art methods in terms of action clarity, especially of the hand motion in the target region, across diverse environments and actions.  [^1] Video results can be found on the Project Page
@@ 1 Introduction
Videos have become commonplace to learn the intricacies of new skills or refresh forgotten ones, such as cooking a certain dish or fixing a broken bicycle [88, 9, 81, 82, 98, 4, 43, 3, 80, 11], including significantly complex skills like surgery [18]. Learning from videos is so central to modern society that more than fifty percent of YouTube viewers seek instructional content [76]. One particularly important class of videos are ones that demonstrate sequences of goal-oriented actions usually involving dextrous hand motion and interaction with tools, objects, and the environment [112, 26, 32, 89]. This class of videos has even demonstrated extensive effectiveness in robot learning [101, 45, 110, 30, 100, 68, 100, 23, 60, 42, 17, 5].
The vision community has generated numerous relevant video benchmarks, like Assembly101 [74], YouCook2 [112], Meccano [70], and HoloAssist [87]. However, these benchmark videos are rarely available in arbitrary working environments where humans or robots operate. Although humans possess varying abilities to learn under environment variation and distraction [24], such factors pose significant challenges for robots or other AI agents, limiting the applicability to broader scenarios [79]. Recent work has explored the generation of instructional images [47, 79] or the retrieval of existing videos for reference [4, 33], with limited success. But, images do not highlight the temporal flow of actions, requiring challenging extrapolation, and the retrieved videos have various distractions, such as a different visual appearance of target objects, different view points, and background variations, which increase cognitive load for human learners [16] and is actually harmful to robot policy learning [17, 42, 60].
To address this gap, we consider a new problem that seeks to generate such goal-oriented hand-centric videos. We refer to this problem as hand-centric video generation or HCVG. Ideally, the generated videos would be usable for such human or robot skill learning, or they may be generally useful as in the broader case of video generation. Concretely, HCVG assumes text and image input, similar to image generation work [47, 79]. The text describes the target instructional action and the RGB image provides the context of the visual environment. As shown in Fig. 1, the generated hand-centric videos are expected to showcase the consistent motion of the target action with a clear subject (the hand and any tools and objects in use) appearance.
The intricacies of this complex hand-centric motion are easily observed in any of the popular instructional video datasets, like HowTo100M [62], Recipe1M [61], and YouCook2 [112]. At the same time, the generated videos should not introduce any unnecessary distractions: e.g., after “cutting a carrot,” the carrot should actually be rendered as a bunch of carrot pieces, but the rest of the environment should not be changed (no hallucinations in the background, no lighting changes, no camera movements, and so on).
Importantly, the new HCVG problem is different than the instruction-based editing problem in works like InstructPix2Pix for images [12], 3D [35, 95], and video [15, 31, 106]. Although many of these works use an image-conditioned diffusion model, the tasks and necessary modeling are distinct. In hand-centric video generation, given an image and a text prompt, the task is to generate a video given the desiderata described above. Likewise, HCVG is different than the numerous text-to-video works [107, 59, 53, 40, 91], as these works do not support an image as input context with the text prompt, and such an extension is non-trivial.
On the other hand, HCVG is closely related to the well-known text-and-image to video (TI2V) problem. TI2V methods, such as TI2V-Zero [66] and PIA [109], extend text-to-image (T2I) models [71, 72] to produce videos conditioned on images and text. However, there are critical differences between them that, despite the advances in TI2V, make HCVG a distinct challenge. As we detail below, these include the ability to avoid distracting content and to generate high-quality hand motions with realistic appearance including object effects from the actions.
First, existing TI2V work focuses on content that is scenery-, character- or object-centric. In those videos, the motion area (MA)—the spatial region in the input image (and generated video) in which to focus the generation of target actions; MA is a term we adopt from AnimateAnything [25]—dominates because of the large target size and the typically clean background (for character- and object-centric cases). In contrast, the hand-centric videos in study usually have fine-grained motion with a small subject (e.g., hand) carrying out a well-defined task, surrounded by arbitrary, sometimes quite cluttered backgrounds. Unfortunately, as shown in Fig. 1 contemporary TI2V methods lead to distracting content like a hallucination in the background (e.g., a hand pops up).
Even though some work has identified these challenges in hard samples, they rely on the extra user input of the MA [25] or even motion trajectory [105, 52], which does not scale. Moreover, the MA is a function of the specific activity and may vary throughout the generated video; hence in HCVG it must be automatically computed. Recent works about robotic policy learning via video generation [27, 101, 78] mitigate the hallucination issue by fitting a huge amount of training samples, but they still struggle with camera movement in the generated videos because the training videos contain significant amounts of it, which is known to be highly distracting for learners [13, 44].
Second, generating high-quality articulated hand motion is important but hard. The generated hand must be consistent—smooth and non-flickering hand pose in consecutive frames. And, yet, consistent is not enough: it must also capture the semantics of the task described by the prompt, showcasing the correct motion and interaction with the environment, tools, and objects within it, as depicted in the input image and necessitated by the text prompt. Existing visual generation techniques, such as VideoCrafter [21, 20] and DragNUWA [105] among many, e.g.,  [67, 64, 57, 102], focus on general-purpose synthesis but are less effective in meeting these critical HCVG needs.
To overcome these technical challenges, we propose a novel hand-centric video generation method—we call it HANDI for hand-centric diffusion—that automatically infers the MA and incorporates a Hand Refinement Loss with a 3D diffusion model (space-time) as the backbone [39, 71]. HANDI has a two-stage approach for HCVG, both stages are diffusion-based in 3D space-time. In the first stage, we learn a model to infer the MA from the input image and text prompt. In the second stage, we augment the diffusion backbone with an innovative Hand Refinement Loss, since the consistency of motion and appearance of the hand is critical in HCVG. Importantly, we incorporate the learned motion area mask so that the stage 2 diffusion model can implicitly learn about object state changes that occur when the hand is carrying out the action specified in the prompt, without a need to explicitly model the object state.
@@
We augment two previous action-oriented image generation datasets [47] based on Ego4D [32] and EpicKitchens [26] for evaluation. In addition to various standard video generation metrics, we propose two additional metrics focusing on good MA and hand appearance. We first compare with SOTA video generation methods to show HANDI’s effectiveness, demonstrating our method’s strong performance for HCVG. Then, we take a closer look at model performance along key axes: (1) the performance of samples with large MA and motion complexity defined by hand movement, (2) the ability to capture the fine-grained motion of the hands, (3) the performance within the MA versus the whole frame, and (4) compute time needed given the two-stage nature of our method. We show that our model outperforms SOTA baselines along each of these axes. Our key contributions are two-fold:
@@
    We demonstrate that HANDI is able to outperform state of the art video generation from image and text-prompt methods, along all measurement axes.
All code, data, and benchmarks will be open-sourced, including the ability to reproduce every included result.
@@ 3 Experiments
Datasets Existing video generation benchmarks mostly feature aesthetically filtered or professionally edited web videos, covering diverse topics [8, 10, 41, 50, 72, 86, 2, 73, 7, 22, 96]. We instead evaluate our method on two widely used hand-centric interaction video benchmarks: Epic-Kitchens-100 (EK) [26] and Ego4D [32], using the quality-checked versions released by LEGO [47]. EK contains 60429/8857 train/test samples and Ego4D contains 85521/9931 train/test samples. The average video length are 2.4/1.07 seconds for EK/Ego4D. We choose EK and Ego4D as they are raw and unfiltered recordings focusing on hand actions, manually captured by research teams. As a result, they reflect real-world challenges such as cluttered backgrounds, weak lighting conditions, and natural scene variations while maintaining diversity in actions (97/1772 unique verbs in EK/Ego4D), objects (300/4336 unique nouns in EK/Ego4d), and environments (45/74 kitchens/locations in EK/Ego4D).
Method Parameters Inf. Time (s)
LFDM [65] 0.108 3.57
AA [25] 1.837 5.74
AVDC [104] 0.229 66.23
PIA [109] 1.483 13.8
Open Sora [111] 1.484 44
DynamiCrafter [93] 1.876 9.32
CogVideoX [103] 4.978 108.41
HANDI 3.674 8.64
Table 1: Efficiency results for the HANDI method on EpicKitchens benchmark. The table lists various parameters and their corresponding efficiency values.
@@ 3.1 Comparisons with the State-of-the-Art
Quantitative and Qualitative Results Table 4 details our compelling quantitative comparisons. HANDI performs best across nearly all metrics and SOTA baselines, including the motion intensive samples—the subset of EK that have top 10% largest $M_{video}$2.1). In a few cases (4 values among 24), strong baselines (AVDC and AA) perform best; yet, neither shows consistent advantage across any metrics on all benchmarks. For example, although AVDC achieves the lowest FVD in EpicKitchens (the full set and motion-intensive subsets), it ranks the third for FVD on Ego4D, which is a more challenging benchmark with more diverse actions, objects, and environments (i.e., outside of the kitchen [32]). Notably, HANDI also produces the most semantically relevant videos, reflected in the highest $\text{CLIP}_{Tx.}$ and BLIP scores, which measure alignment with the text prompt. We discuss the (strong performance in the) hand structure error below in §3.2.
@@
Computational Efficiency Analysis Given HANDI’s two-stage design, one consideration is computational efficiency. Despite introducing an additional stage for motion area generation, our method remains highly efficient as it allows us to use a lightweight model structure for both stages to handle the easier MA task. We compare inference times for generating one video across the latest baselines: DynamiCrafter requires 9.32 seconds, CogVideoX takes 108.41 seconds, while HANDI completes generation in just 8.64 seconds, demonstrating that HANDI not only achieves superior video quality but also operates efficiently.
@@ 3.2 Ablation Study
We conduct an ablation study to investigate the effect of our two new model elements—generated Motion Area (MA) and Hand Refinement Loss (HRL); shown in Tab. 3 and Fig. 5. In the ablation, we not only assess performance on the full frames (Scope is “Full”) but also assess it on the ground truth motion area $M_{video}$2.1) since this is the detailed hand-centric part of the video we focus on in HCGV (Scope is “MA”). Note that no mask (Mask is $\times$) means the motion area is set to the full frame for stage 2 generation.
@@ 4 Conclusion
We tackle the unique challenges in generating videos of hands performing goal oriented actions, an understudied aspect of video generation that is important for numerous applications, such as human and robot learning. In these hand-centric video generation tasks, one seeks high-fidelity hand structure with a stable video background based on the provided image and text conditioning. Our method, called HANDI, builds on the state of the art diffusion-based video generation literature with two important innovations. First, we automatically generate the motion area based on the visual context and the action prompt, allowing HANDI to focus on the region in the scene that is most relevant to the conditioning. Second, we propose a key hand refinement loss term that drives the diffusion model to focus on learning to generate high quality, realistic hand poses. Our SOTA comparisons and baselines on challenging real-world video datasets uniformly and significantly show the power of these two innovations. We have identified new avenues of future work for video generation in realistic hand-centric settings; for example, we plan to consider explicit models of the objects and tools in use for future work.
Acknowledgement
We would like to thank the anonymous reviewers for their valuable comments and suggestions. This work is sponsored by Defense Advanced Research Projects Agency (DARPA) HR00112220003 (Perceptually-enabled Task Guidance). The contents do not reflect the position of the U.S. Government.
Appendix A Ablation Study Qualitative Results Shown in Videos
For the best visualization of motion, we encourage readers to visit the videos for ablation study qualitative results, attached ablation.mp4 or in project page.
Appendix B More Qualitative Results Comparing with SOTA methods Shown in Videos
Figure 6: An overview of the HTML page that shows more qualitative results in videos. Please view the full page in attachment baseline_comparison/index1.html or the online project page.
Likewise, for comparison with SOTA methods, we encourage readers to view 10 qualitative results in videos in the off-line HTML page attached  baseline_comparison/index1.html or the online project page. A screenshot of the page is shown in Fig. 6 for readers’ reference.
By comparing with all baselines, we emphasize three key capabilities of our method that are accountable for the two key designs–Motion Area (MA) Generation and Hand Refinement Loss (HRL). Specifically, we observe good performance on generating Hand-Centric videos for actions require (i) intricate fingertip motion; (ii) large Motion Area; (iii) Object State Changing. Our proposed learnable automatic Motion Aera mask generation helps the model to focus on the accurate region for the action execution (“ii” and “iii”), avoiding adding distraction to background. Our novel Hand Refinement Loss helps the model handle the actions with critical details (“i”), which is under-explored in previousw ork. In this gallery, for each sample, the first row displays the input context image and the results from different methods. In the second row, we first present the predicted Motion Area mask from our Stage 1 model. Then in the same row, we show the motion flow below generated video from each method. Note that the motion flow is not part of the model’s outputs; it is shown solely to facilitate easier comparison among methods. Below the visuals, we provide the input Text Prompt for the target action (e.g., ”Pick up and crack egg.”) and the Enriched Action Description (e.g., ”The person uses their left hand to pick up an egg from the egg box and cracks it into a bowl.”).
Appendix C BLIP Score with Different BLIP Model Size
@@
Appendix D Generation of Various Actions in the Same Visual Context
We explore our model’s performance on generating different actions given the same input context image. These are novel image-action pairs that have never shown in the dataset including the test set. We illustrate one example in Fig. 7 that crosses two existing samples in the test set. These results further indicate that our model can generate Hand-Centric videos for various actions, emphasizing hand movement and subtle fingertip motion.
Figure 7: An illustration that crosses the input action text and context image in two test samples. Our method generates reasonable Hand-Centric videos for such challenging scenarios of subtle actions.
Appendix E Motion Flow
As we shown in the qualitative results in both main paper and supplementary materials, we compute Motion Flow to assist the illustration of the motion in the generated video from each method. We first compute optical flow between all two consecutive frames in the generated video. Then we overlay all optical flows in a way that counting the number of frames that has optical flow intensity for each pixel. Therefore, the produced map has the same size as the optical flow (and the generated video) where each pixel has the intensity $m^{\prime}\in[0,\dots,L]$ (inclusive) where $L$ is the number of generated frames. In the end, we normalize the intensities by $L$ to produce our visualized Motion Flow where each pixel has the intensity $m\in[0,1]$ (inclusive).
Appendix F Baseline Implementation Details
Since previous Text-Image-to-Video/Image Animation work does not report results on the Hand-Centric Video Generation (HCVG) task, we implement all four baselines. For a fair comparison, we train or fine-tune their methods on the same training set used for our own method. For DynamiCrafter [93], CogVideoX [103], and Open Sora [111], we directly run the published codebase with fine-tuning and inference code to obtain our reported results. We provide more details about the implementation as follows:
F.1 Animate Anything
The Animate Anything [25] model is initialized from VideoComposer [90], which is pre-trained on the WebVid10M [6] dataset, a large-scale video dataset designed for general video generation tasks. For fine-tuning, the model is trained on 20,000 videos randomly sampled from the HDVILA-100M dataset [97]. These samples are carefully selected to remove watermarks, ensuring higher-quality training data. The model leverages a latent diffusion framework to generate fine-grained animations guided by motion area masks and motion strength.
F.2 Latent Flow Diffusion Model (LFDM)
LFDM [65] is a versatile framework with multiple pre-trained versions. The available pre-training datasets include:
  • The MUG Facial Expression Dataset [1]: Contains 1,009 videos of 52 subjects performing 7 different facial expressions.
  • The MHAD Human Action Dataset [19]: Comprises 861 videos of 27 actions performed by 8 subjects.
  • The NATOPS Aircraft Handling Signal Dataset [77]: Consists of 9,600 videos of 20 subjects performing 24 body-and-hand gestures used for communicating with U.S. Navy pilots.
For the experiments, the version pre-trained on the MHAD Human Action Dataset is used. LFDM generates 8-frame video sequences, employing a two-stage training process: an unsupervised latent flow auto-encoder and a conditional 3D U-Net-based diffusion model for temporal latent flow generation. Training is computationally efficient, requiring only 4 GPUs for a single day of training.
F.3 Personalized Image Animator (PIA)
PIA is trained on the WebVid10M dataset [6], focusing only on the condition module and temporal alignment layers as trainable parameters. The motion module from AnimateDiff [34] is used as a pre-trained model for initializing the temporal alignment layers. PIA fine-tunes on curated images generated by personalized text-to-image (T2I) models, combined with motion-related prompts. These curated images and prompts are included in the AnimateBench benchmark, which consists of diverse images across multiple styles and tailored motion-related annotations. Fine-tuning focuses on the condition module and temporal alignment layers while freezing the base T2I model.
F.4 Actions from Video Dense Correspondences (AVDC)
AVDC [104, 46] is trained on a collection of video datasets featuring table-top manipulation and navigation tasks. Although the specific dataset names are not provided, the training framework synthesizes 8-frame video sequences using a text-conditioned video diffusion model. The weights of the ResNet-18 [36] backbone are initialized either from scratch (BC-Scratch) or using the pre-trained parameters of R3M [63]. GMFlow [94] is employed for optical flow prediction, enabling dense correspondences for action regression. The model training is computationally optimized to run on 4 GPUs within a single day.

### v5

@@ Abstract
Despite the recent strides in video generation, state-of-the-art methods still struggle with elements of visual detail. One particularly challenging case is the class of videos in which the intricate motion of the hand coupled with a mostly stable and otherwise distracting environment is necessary to convey the execution of some complex action and its effects. To address these challenges, we introduce a new method for video generation that focuses on hand-centric actions. Our diffusion-based method incorporates two distinct innovations. First, we propose an automatic method to generate the motion area—the region in the video in which the detailed activities occur—guided by both the visual context and the action text prompt, rather than assuming this region can be provided manually as is now commonplace. Second, we introduce a critical Hand Refinement Loss to guide the diffusion model to focus on smooth and consistent hand poses. We evaluate our method on challenging augmented datasets based on EpicKitchens and Ego4D, demonstrating significant improvements over state-of-the-art methods in terms of action clarity, especially of the hand motion in the target region, across diverse environments and actions.
@@ 1 Introduction
Instructional media have become the de-facto manual for everything from “how to poach an egg” videos to full laparoscopic surgery walkthroughs [61, 11]. More than half of YouTube users now search specifically for how-to content [52]. Within this vast stream, the clips most useful for both humans and robots are those that capture goal-oriented hand manipulation: fine, dexterous interactions between hands, tools, and objects performed in ordinary, cluttered environments [81, 15, 19, 62]. Such footage accelerates human skill acquisition and has proved invaluable for imitation-based policy learning [72, 27, 10].
Obtaining task-specific videos on demand remains difficult. Retrieval systems often return clips recorded in different kitchens, workshops, or lighting, breaking visual continuity with the user’s scene [2, 20]. Single-frame generators can preserve context but sacrifice temporal dynamics [28, 55]. Recent text-and-image-to-video (TI2V) models attempt to offer both context and motion [43, 79], yet they are designed for web benchmarks featuring centred subjects against clean backgrounds [4, 5, 24, 31, 48, 59, 1, 49, 3, 13, 70]; when confronted with real household footage they hallucinate background motion and render low-fidelity hands, especially when the hands occupy only a small portion of the frame (see Fig. 1). Bridging this gap is essential for practical human-assistive and robot-assistive scenarios.
Motivated by these limitations, we cast the task as hand-centric video generation (HCVG). Given (i) a single RGB scene image that shows the workspace before execution and (ii) a text command describing the desired manipulation, HCVG requires synthesising a video that (a) alters pixels only inside the motion area—the small, task-specific region encompassing the hand, tool, and target object—while leaving the surrounding context untouched; (b) produces anatomically correct, temporally smooth hand motion; and (c) realises any downstream object-state change required by the action (e.g., transforming a whole carrot into diced pieces after cut carrot; see  Fig. 1). In another word, HCVG emphasize the same input and output modality as TI2V but focuses on animating high-quality hands from challenging, unedited real-world images. Note that, unlike instruction-based TI2V, which enjoys more information (e.g., the motion area or action trajectory) from the user [6, 14, 76], HCVG requires less user burden and requires the generator itself to localise and track this region in cluttered scenes where the hand often occupies less than a quarter of the frame.
To meet these demands we introduce HANd-centric DIffusion (HANDI), a two-stage 3-D diffusion framework. Stage 1 automatically predicts a spatio-temporal motion-area (MA) mask from the image–text pair, using a lightweight video-diffusion backbone . We develop a pipeline to automatically prepare the pseudo MA masks from training videos for training. Stage 2 conditions the backbone [23, 47] on the predicted mask to generate the instructional video clip. This stage optimises a Hand-Refinement Loss for hand-pose consistency. HRL supervises by comparing the generated and groundtruth hand skeletons, across all generated frames, introducing explicit hand structure modeling among plain pixels. This two-stage MA-aware generation framework and HRL allow HANDI to render high-fidelity hands even in the presence of severe clutter.
@@
To benchmark HCVG, we use two large-scale hand action datasets, Ego4D [19] and EpicKitchens [15], featuring various unedited real-world visual backgrounds. In addition to various standard video generation metrics, we propose two additional metrics focusing on good MA and hand appearance. We first compare with SOTA video generation methods to show HANDI’s effectiveness, demonstrating our method’s strong performance for HCVG. Then, we take a closer look at model performance along key axes: (1) the performance of samples with large MA and motion complexity defined by hand movement, (2) the ability to capture the fine-grained motion of the hands, (3) the performance within the MA versus the whole frame, and (4) compute time needed given the two-stage nature of our method. We show that our model outperforms SOTA baselines while maintaining computation efficiency.
Our key contributions are two-fold:
@@
    We demonstrate that HANDI is able to outperform state of the art video generation from image and text-prompt methods, along all measurement axes
2 Related Work
Position of our work. HANDI unifies the desirable properties of several research threads while removing their respective limitations. Unlike image generation work [28, 55], our method produces dynamics—temporally coherent demonstrations. Unlike text-only video models (I2V), which ignore the user’s scene and therefore hallucinate context [77, 65], HANDI conditions on the real workspace image and keeps the visual context like object appearance matched. In contrast to prompt-style Text-Image-to-Video (IT2V) generation work that require masks or trajectories from the user [14, 76], HANDI does not require those scribbles and lowers user burden, enabling scalable deployment in cluttered, real-world environments where the hand occupies only a tiny fraction of the view. Finally, compared with standard TI2V systems that animate the whole frame and blur small manipulators [43, 79], we explicitly localise the motion area and refine 3-D hand pose, yielding the right amount of high-fidelity articulation.
Instructional-video corpora. Large egocentric datasets such as Assembly101 [50], YouCook2 [81], Meccano [46], HoloAssist [60], Ego4D [19] and EPIC-Kitchens [15] capture sequences of human-object interaction and drive progress in action recognition and imitation learning. Their footage is unedited: hands are small and backgrounds cluttered.
Image- and video-based visual instruction. Context-preservation is the key to providing instruction. Viewpoint and appearance mismatches increase cognitive load for humans [9] and degrade robot policies [38]. Generating static, context-aware instructional images has been explored by LEGO [28] and GenHowTo [55], but images lack temporal cues essential for manipulation. Video generation conditioned on text alone has advanced rapidly [77, 65], yet these models produce videos with mismatched context. Text-and-image-to-video (TI2V) systems such as TI2V-Zero [43] and PIA [79] add scene context but still treat the whole frame as a canvas, resulting in background motion and distorted hands. Recent instruction-based editors (e.g., Pix2Video [8], AnimateAnything [14], DragNUWA [76]) mitigate distractions by requiring user-supplied masks or keypoint trajectories, limiting scalability.
Motion-area localisation and hand fidelity. When the manipulation locus is tiny—often $<\!25\%$ of the frame—automatic localisation becomes critical. Prior work either relies on large-scale domain randomisation [16, 54] or manual markup [14, 76] and still exhibits camera drift. General-purpose video diffusers such as VideoCrafter [12] or latent transformers [44] improve realism but do not model articulated hand structure, leading to flicker and anatomical errors. Hand-specific refiners for images (HandDiffuser [41], HandRefiner [36]) do not address temporal coherence.
@@ 4 Experiments
Datasets We evaluate on two raw hand-action videos corpora: Epic-Kitchens-100 (EK) [15] and Ego4D [19], using LEGO’s quality-checked splits [28]. EK has 60,429/8,857 train/test samples (2.4s avg.), and Ego4D has 85,521/9,931 (1.07s avg.). These datasets reflect real-world challenges: clutter, lighting variation, and small hand regions—spanning 97/1772 verbs, 300/4336 nouns, and 45/74 kitchen or location settings in EK/Ego4D.
@@ 4.1 Comparisons with the State-of-the-Art
Quantitative and Qualitative Results. Table 3 details our compelling quantitative comparisons. HANDI performs best across nearly all metrics and SOTA baselines, including the motion intensive samples—the subset of EK that have top 10% largest $M_{video}$3.1). In a few cases (4 values among 24), strong baselines (AVDC and AA) perform best; yet, neither shows consistent advantage across any metrics on all benchmarks. For example, although AVDC achieves the lowest FVD in EpicKitchens (the full set and motion-intensive subsets), it ranks the third for FVD on Ego4D, which is a more challenging benchmark with more diverse actions, objects, and environments (i.e., outside of the kitchen [19]). Notably, HANDI also produces the most semantically relevant videos, reflected in the highest $\text{CLIP}_{Tx.}$ and BLIP scores, which measure alignment with the text prompt. We discuss the (strong performance in the) hand structure error below in §4.2.
@@
Computational Efficiency Analysis. Although HANDI introduces a separate stage for motion–area prediction, its overall run-time is on par with (or better than) comparable single-stage systems. On EpicKitchens, synthesising a $256\times 256$ clip on one NVIDIA H100 takes 8.6 s. For reference, the light but less expressive AnimateAnything completes in 5.7 s [14], DynamiCrafter in 9.3 s [68], PIA in 13.8 s [79], OpenSora in 44 s [80], AVDC in 66 s [75], and CogVideoX in 108 s [74]. HANDI therefore delivers its hand-fidelity gains without imposing a meaningful latency penalty.
@@ 4.2 Ablation Study
We conduct an ablation study to investigate the effect of our two new model elements—generated Motion Area (MA) and Hand Refinement Loss (HRL); shown in Tab. 2 and Fig. 5. In the ablation, we not only assess performance on the full frames (Scope is “Full”) but also assess it on the ground truth motion area $M_{video}$3.1) since this is the detailed hand-centric part of the video we focus on in HCGV (Scope is “MA”). Note that no mask (Mask is $\times$) means the motion area is set to the full frame for stage 2 generation.
@@ 5 Conclusion
We tackle the challenges in generating videos of hands performing goal-oriented actions, an understudied but important aspect of video generation for applications like human and robot learning. These scenarios demand high-fidelity hand motion, conditioned on a single image and text prompt. Our method, HANDI, builds on diffusion-based video generation with two innovations: automatic localisation of the motion area, avoiding background hallucination, and a hand refinement loss that encourages anatomically realistic and temporal consistent hand motion. Our comparisons on real-world datasets demonstrate consistent gains over prior methods while maintaining efficiency.
Appendix A BLIP Score with Different BLIP Model Size
@@
Appendix B Generation of Various Actions in the Same Visual Context
We explore our model’s performance on generating different actions given the same input context image. These are novel image-action pairs that have never shown in the dataset including the test set. We illustrate one example in LABEL:fig:various_prompt that crosses two existing samples in the test set. These results further indicate that our model can generate Hand-Centric videos for various actions, emphasizing hand movement and subtle fingertip motion.